APACHE POINT OBSERVATORY SDSS 2.5M OBSERVING LOG Tuesday January 8, 2008 (MJD 54474) ---=== OBSERVING TEAM ===--- Kaike Pan (observer) Elena Malanushenko (observer) ---=== OBSERVING PLAN ===--- Science! ---=== OBSERVING SUMMARY ===--- Cirrus for whole night. We observed spectra. We had GALIL problem that caused the secodary fail to move. Rebootting of Galil and homing the secondary fixed the problem. Obtained good signal on plates 1878,2608, and 2640 enough to declare them sa "DONE". Obtained some signal on other plates 2866,1636,2867. ---=== OBSERVING LOG ===--- Afternoon: ----------- Everything was checked out and no problem to report. Spectro focus was checked and no adjustment made, still set for about 1 C. Cart 1 was mounted for the begining of the night. ad3 display does not work again. Night: ------ 00:30 begin ventilation 01:20Z enclosure is off. 00:54Z Cart 1, plate 2866, segueFaint, pri 4. Used FK5, but need to apply large offset to find it. After getting guiding stars, Kaike applied a scale change of 0.01 that resulted lost stars. Then the telescope was sent back to FK5, and found the star was still there as a big donut. Adjusting the focus by +/- 300 did not change the image. Then we back to the field, A careful examining showed guiding stars (as very faint donuts) were actually there. applying a scale change of -0.01 got better images. At the point, we realized that someting happen to the seconary which did not move properly or did not move at all. Thanks to French being on site, we had him to examine the problem. French let us to home/relax the secondary, but we could not get prompt back after HOST TCC_SEC command. Then we went to the telescope, French check the sendary control box, and M2 Galil, they were looked okay. French thought that there was a communication problem. So he decided to power cycle the TPM M2 galil, and had us to home/relax the secodary. This fixed the problem. About 40 min was lost to the problem. 2:11Z goToField again. Used FK5 to find GS. 2:07Z preCalib 2:34Z science 1500 3:02Z science 1500 Plate is expired at 3:24Z. 3:30Z postCalib We obtained good signal on this plate but still need more observations. 3:38Z Cart 7, plate 2640, baoTest, pri 1. FS on GTF. Seeing is below 1". 3:47Z preCalib 3:53Z science 1200 4:16Z science 1200 sec -> 900 sec 4:34Z science 900 sec --> 1500 sec We used the criteria r1,r2 > 20 and expected the plate is done after the last exposure. After we replaced it to another plate already we found out that (s/n)^2 for r1 ~ 19.4 and for r2 ~ 19.3. We declared this plate as "DONE" anyway. 5:08Z Cart 2, plate 1636, photoz. 05:17Z preCalib 05:27Z science 1500 sec exposure 05:55Z science 1500 sec, exposure is aborted because of low efficiency of signal during the exposure, clouds. We obtained some signal on this plate and need to continue its observations. 06:27Z Cart 3,plate 1878, chunk with pri 1. FSonGTF. 03:39Z preCalib 06:46Z science 1200 07:09Z science 1200 --> 1500 07:37Z science 1500 08:05Z postCalib This plate is DONE with (s/n)^2 15.8. 08:14Z Cart 8, plate 2867 FSonGTF 08:30Z preCalib 08:37Z science 1500 09:05Z science 1200 09:28Z postcal We obtained only s/n ~ 9, and need to continue observations. 9:42Z Cart 6, plate 2608, chunk This plate has previous fotons (s/n)^2 =9.7. Used FK5 to find the guider stars. Seeing is very variable in range 1.3-2.5". 09:59Z preCalib 10:06Z science 900--> 1200 10:29Z science 1200-->900 We obtained (s/n)^2 8.5 and this plate is DONE. 10:50Z Cart 4 plate 2609 FSonGTF 11:03 preCalib seeing is 2-3" based on DIMM. 11:09Z science 1500 11:37Z science 1200 --> 900 and endStare with exposure 600 sec because of high humidity. We ignored second exposure and did not include its signal do database because the duration of second exposure is too small compare the first exposure. This plate needs more observations. Humidity is up to 90% and difference dropped to 2.0C. 11:55Z Closed 12:10Z endNight is completed. ---=== IMAGING RUN SUMMARY ===--- Run Time Stripe Lambda Last Flavor Comments Start End Begin End Frame ------------------------------------------------------------------------- 7274 22:31Z 22:49Z 100 O 154.72 159.23 44 ignore ---=== IMAGING RUN DETAILS ===--- ---=== SKIPPY RESULTS ===--- Run Frame nFrames stars muErr muRms nuErr nuRms rot az el --------------------------------------------------------------------------- ---=== LTMATCH RESULTS ===--- Run Field nFields alt az nGood rowMean rowSig colMean colSig rot ------------------------------------------------------------------------ ---=== SPECTROSCOPY DATA SUMMARY ===--- Summary Checked (y/n): yes QA Procedures Done (y/n): yes UT Exp Time flavor comment (S/N)^2 totals ========================================== b1 r1 b2 r2 22:48 48697 3.0 arc ----- sequence 48698, plate -9999 ------- 00:13 48698 0.0 bias ----- sequence 48699, plate 2866 ------- 02:29 48699 10.0 flat 02:32 48700 2.0 arc 03:01 48701 1501.0 target 03:29 48702 1501.0 target 03:32 48703 10.0 flat 03:35 48704 2.0 arc ----- sequence 48705, plate 2640 ------- 03:50 48705 10.0 flat 03:52 48706 2.0 arc 04:16 48707 1201.0 target 04:34 48708 901.0 target 05:02 48709 1502.0 target ----- sequence 48710, plate 1636 ------- 05:20 48710 10.0 flat 05:22 48711 2.0 arc 05:55 48712 1501.0 target ----- sequence 48713, plate 1878 ------- 06:42 48713 10.0 flat 06:44 48714 2.0 arc 07:08 48715 1200.0 target 07:36 48716 1503.0 target 08:04 48717 1500.0 target 08:07 48718 10.0 flat 08:10 48719 2.0 arc ----- sequence 48720, plate 2867 ------- 08:32 48720 10.0 flat 08:35 48721 2.0 arc 09:04 48722 1501.0 target 09:27 48723 1201.0 target 09:30 48724 10.0 flat 09:33 48725 2.0 arc ----- sequence 48726, plate 2608 ------- 10:02 48726 10.0 flat 10:05 48727 2.0 arc 10:28 48728 1202.0 target 10:46 48729 903.0 target ----- sequence 48730, plate 2609 ------- 11:06 48730 10.0 flat 11:08 48731 2.0 arc 11:37 48732 1501.0 target 11:50 48733 628.0 target ---=== TELESCOPE OFFSETS AND SCALE I ===--- Time Instrument Az Alt Rot Scale pos offset pos offset pos offset ------------------------------------------------------------------------------ 02:42Z 1 2866 -22.65 0.0058 55.14 0.0057 161.10 0.0148 1.000110 04:27Z 7 2640 -24.12 0.0040 54.84 0.0040 159.91 0.0120 0.999910 06:54Z 3 1878 158.73 0.0057 49.10 0.0036 -51.96 0.0081 1.000170 09:58Z 6 2608 70.22 0.0074 61.33 0.0022 237.20 0.0081 1.000160 12:28Z 4 2609 121.00 0.0074 29.87 0.0022 0.00 0.0081 1.000140 ---=== TELESCOPE OFFSETS AND SCALE II ===--- ---=== FOCUS LOG ===--- setmir piston Temp Wind Time Inst scale M1 M2 Foc Az Alt (C) MPH Dir filt fwhm ------------------------------------------------------------------------------ 02:42Z 1 2866 1.00011 -1086 336 70 -23 55.1 1.5 4 257 -- 1.2" 04:27Z 7 2640 0.99991 889 1952 50 -24 54.8 2.5 6 267 -- 1.2" 06:54Z 3 1878 1.00017 -1679 -156 50 159 49.1 2.2 14 250 -- 1.7" 09:58Z 6 2608 1.00016 -1580 142 300 70 61.4 -0.3 12 281 -- 1.5" 12:28Z 4 2609 1.00014 -1383 233 100 121 29.9 -2.0 12 268 -- 1.0" ---=== WEATHER LOG ===--- Wind Time Temp F Dewp F MPH Direction Dust DIMM Sky 22:47Z 41 -7 5 303 (WNW) 222 - 23:18Z 39 2 0 349 (NNW) 246 - 23:52Z 39 -8 2 280 (W) 136 - 00:24Z 35 -10 5 270 (W) 180 - 00:56Z 37 -11 2 203 (SSW) 123 - 01:28Z 31 0 7 253 (WSW) 238 - 01:58Z 34 -14 7 248 (WSW) 160 - 02:29Z 34 -13 10 266 (W) 157 - 03:00Z 33 -15 6 259 (W) 173 - 03:31Z 34 -14 11 269 (W) 147 - 04:01Z 35 -13 10 267 (W) 143 - 04:33Z 36 -13 9 267 (W) 187 - 04:57Z 36 -13 10 268 (W) 137 - 05:35Z 36 -13 11 268 (W) 164 - 06:06Z 36 -13 11 265 (W) 180 - 06:38Z 35 -14 11 270 (W) 189 - 07:08Z 34 -14 7 220 (SW) 173 - 07:38Z 30 -11 11 282 (WNW) 161 - 08:10Z 30 4 14 281 (W) 168 - 08:41Z 30 18 16 290 (WNW) 470 - 09:11Z 30 18 16 293 (WNW) 448 - 09:43Z 30 20 8 263 (W) 489 - 10:13Z 30 20 14 272 (W) 443 - 10:44Z 31 20 15 260 (W) 432 - 11:14Z 30 22 10 269 (W) 517 - 11:45Z 29 25 11 263 (W) 632 - 12:16Z 28 26 10 266 (W) 807 - ---=== TELESCOPE STATUS ===--- 00:30Z - Began ventilation. Doors and louvers opened, fans on. 01:20Z - Enclosure off. 12:00Z - Enclosure on, doors and louvers closed, fans off. No interlocks bypassed. Status at: 12:30Z Telescope stowed at: 30,121,00 mount Instrument mounted: Cartridge 4 Counterweights at: 270 lbs LN2 autofill systems: Connected and turned on. 180L LN2 dewar scales: Spectro 256 lbs, 18 psi Imager 167 lbs, 22 psi ---=== SOFTWARE USED ===--- IOP/SOP: v4_38_0 Watcher: v2_36_0 MCP: v6_4_0 TPM: tpm_v3_5_0 AstroDa: v15_13_1 TCC: 1819700 packets sent successfully sdssProcedures: v2_37 SoS: v5_1_8 plate-mapper: v4_3_1 ---=== MIRROR NUMBERS ===--- PRIMARY: -------- Scale: 1.000000 MIGS TONIGHT NOMINAL Axial A -5.2160 -5.2160 Axial B -0.2142 -0.2142 Axial C -0.0015 -0.0016 Trans D 11.8440 11.8540 Lateral E 10.0711 10.0584 Lateral F 11.6586 11.6586 GALILS Commanded: 4600. -2000. 2650. -12450. -6500. -6550. Actual: 4600. -2000. 2650. -12439. -6477. -6546. SETMIR VALUES PriDesOrient: 0.00 0.00 0.00 100.00 -200.00 PriOrient: 0.00 0.00 0.00 99.59 -200.69 SECONDARY: ---------- Focus: 0.00 Air Temp.: 5.1 Alt.: 29.999993 MIGS TONIGHT NOMINAL Axial A 0.8027 0.8025 Axial B 0.6122 0.6120 Axial C 0.5474 0.5474 Trans D 0.6005 0.6008 GALILS Commanded: 1627586. 1652631. 1470887. -5900. -7800. Actual: 1627496. 1653168. 1470885. -5887. -7780. SETMIR VALUES SecDesOrient: 1257.00 60.00 -4.00 -20.00 118.90 SecOrient: 1256.88 59.88 -4.17 -20.03 119.36 ---=== PROBLEMS IN DETAIL ===--- sp1> mechSend s Coll_motor_A -11051 01P-0000011051 Coll_motor_B -11028 02P-0000011028 Coll_motor_C -12137 01P-0000012137 Temp_Median 5.9 sp2> mechSend s Coll_motor_A -3943 01P-0000003943 Coll_motor_B -4889 02P-0000004889 Coll_motor_C -5041 `1P-0000005041 Temp_Median 3.7 IDL 5.5a> collimate, 48696 COLLIMATE: Camera b1 appears to be OUT-OF-FOCUS (|mean| > 0.15 pix) COLLIMATE: Predict blue camera ring movement of -5.4 degrees for spectro-1 COLLIMATE: (if red is already in focus) COLLIMATE: Mean offset = 0.16 pix COLLIMATE: Camera b2 appears to be OUT-OF-FOCUS (|mean| > 0.15 pix) COLLIMATE: Predict blue camera ring movement of -5.7 degrees for spectro-2 COLLIMATE: Mean offset = -0.22 pix COLLIMATE: Camera r1 appears to be OUT-OF-FOCUS (|mean| > 0.15 pix) COLLIMATE: Predict (red) piston movement of 2032 steps for spectro-1 COLLIMATE: Mean offset = -0.28 pix COLLIMATE: Camera r2 appears to be OUT-OF-FOCUS (|mean| > 0.15 pix) COLLIMATE: Predict (red) piston movement of 2598 steps for spectro-2 COLLIMATE: Issue SOP command: "sp2; mechPiston 2598" ----------- Watcher complains: m2 galil (Communications Fault: 2008-01-09 02:05:07Z) ------- Spectro bias: Filename 02% 05% 10% 50% 90% 95% 98% ------------------- ------ ------ ------ ------ ------ ------ ------ sdR-b1-00048698.fit -5.7 -4.1 -2.4 2.2 7.4 8.6 10.6 sdR-b2-00048698.fit -7.8 -6.2 -4.9 1.2 7.4 8.6 11.1 sdR-r1-00048698.fit -7.2 -5.2 -4.1 1.3 6.8 8.8 10.0 sdR-r2-00048698.fit -7.6 -5.5 -4.3 1.0 6.2 8.2 10.3 ----------